# On Limitations of LLM as Annotator for Low Resource Languages

### v1

@@ Abstract
Low-resource languages face significant challenges due to the lack of sufficient linguistic data, resources, and tools for tasks such as supervised learning, annotation, and classification. This shortage hinders the development of accurate models and datasets, making it difficult to perform critical NLP tasks like sentiment analysis or hate speech detection. To bridge this gap, Large Language Models (LLMs) present an opportunity for potential annotators, capable of generating datasets and resources for these underrepresented languages. In this paper, we focus on Marathi, a low-resource language, and evaluate the performance of both closed-source and open-source LLMs as annotators. We assess models such as GPT-4o and Gemini 1.0 Pro, Gemma 2 (2B and 9B), and Llama 3.1 (8B) on classification tasks including sentiment analysis, news classification, and hate speech detection. Our findings reveal that while LLMs excel in annotation tasks for high-resource languages like English, they still fall short when applied to Marathi. Even advanced closed models like Gemini and GPT underperform in comparison to BERT-based baselines, highlighting the limitations of LLMs as annotators for low-resource languages.
@@ 1 Introduction
India is a multilingual nation with various regional languages and most of these languages fall under the low-resource (LR) category because of the scarcity of digital resources.
@@
Specifically, we evaluated models such as GPT-4o, Gemini 1.0 Pro, Gemma 2 (2B and 9B), and Llama 3.1 (8B) across multiple tasks, including 3-class sentiment analysis, 2-class, and 4-class hate speech detection, as well as news classification based on headlines, long paragraphs, and full documents.
The paper is structured as follows: Section 2 provides a concise review of prior research on data annotation and the use of LLMs. In Section 3, we detail the datasets used and the models employed in our evaluation. Section 4 describes the experimental setup and the APIs leveraged to evaluate the LLMs. Section 5 presents the results along with a comparative analysis of various open-source and closed-source LLMs, as well as BERT-based models. Finally, in Section 6, we conclude our discussion.
@@ 2 Literature Review
Bias is another concern with LLMs. Bavaresco et al. (2024) introduced JUDGE-BENCH to evaluate LLM biases, noting that training data heavily influences model outputs, which can be problematic in annotating complex or sensitive tasks in low-resource languages. While LLMs used for high-resource language(HRL) are giving promising results, that is not the case for low-resource languages. Nasution and Onan (2024) explored ChatGPT-4’s performance in annotation tasks across languages like Turkish and Indonesian, offering insights into LLM applicability for Low Resource Language(LRL), a relevant consideration for our focus on Marathi.
| Dataset | Tech | Llama 3.1 8B | Gemma 2 2B | Gemma 2 9B | Gemini 1.0 Pro | GPT-4o | BERT |
| MahaSent | ZS | 0.76 | 0.71 | 0.69 | 0.78 | 0.79 | 0.80 |
| FS | 0.79 | 0.76 | 0.78 | 0.76 | 0.82 | ||
| MahaHate-2C | ZS | 0.64 | 0.71 | 0.78 | 0.74 | 0.80 | 0.91 |
| FS | 0.78 | 0.72 | 0.82 | 0.72 | 0.82 | ||
| MahaHate-4C | ZS | 0.40 | 0.39 | 0.43 | 0.43 | 0.58 | 0.73 |
| FS | 0.48 | 0.41 | 0.46 | 0.45 | 0.60 | ||
| MahaNews-SHC | ZS | 0.60 | 0.54 | 0.68 | 0.68 | 0.78 | 0.85 |
| FS | 0.66 | 0.54 | 0.68 | 0.70 | 0.78 | ||
| MahaNews-LPC | ZS | 0.66 | 0.55 | 0.71 | 0.72 | 0.77 | 0.89 |
| FS | 0.67 | 0.50 | 0.72 | 0.74 | 0.75 | ||
| MahaNews-LDC | ZS | 0.69 | 0.62 | 0.78 | 0.74 | 0.81 | 0.96 |
| FS | 0.69 | 0.62 | 0.80 | 0.75 | 0.81 |
3 Methodology
We investigate the distinctions between LLM-generated and human-generated annotations for the Indic language, Marathi, using a comparative methodology, and analyzed the results with BERT-based models for detailed insights. In this research, we focus on three major task categories using relevant Marathi datasets: 1) MahaSent Kulkarni et al. (2021); Pingle et al. (2023) – classifies sentiment of Marathi tweets into positive, negative, or neutral categories. 2) MahaHate Patil et al. (2022) – measures the level of abusive and hostile content in Marathi text. This dataset includes two supervised tasks: MahaHate 2-Class, which categorizes content as either HATE or NOT, and MahaHate 4-Class, which provides finer distinctions with categories: Hate (HATE), Offensive (OFFN), Profane (PRFN), and Not (NOT). 3) MahaNews Mittal et al. (2023); Mirashi et al. (2024) – classifies headlines and articles from Marathi news. It comprises three supervised datasets: Short Headlines Classification (SHC), Long Document Classification (LDC), and Long Paragraph Classification (LPC), each categorizing news content into 12 classes: Auto, Bhakti, Crime, Education, Fashion, Health, International, Manoranjan, Politics, Sports, Tech, and Travel.
@@ 3.1 LLMs
In our annotation experiments, we evaluated the performance of LLMs for the Marathi language using two prompting techniques: zero-shot and few-shot learning. We tested both open-source models (Llama 3.1 8B, Gemma 2 2B, and Gemma 2 9B) and closed-source models (Gemini 1.0 Pro, GPT-4o), and compared their results with BERT-based models, as detailed in Section 5.1. The performance of each LLM under both prompting strategies is summarized in Table 1.
@@ 3.2 BERT Based Models
We used BERT-based models to compare performance with LLMs, where MahaSent-MD, MahaHate-BERT, MahaNews-SHC-BERT, MahaNews-LPC-BERT, and MahaNews-LDC-BERT are fine-tuned versions of MahaBERT, while MahaHate-multi-RoBERTa has MahaRoBERTa as the base model. Each of these models was fine-tuned on the corresponding datasets, and their respective performances are summarized in Table 1.
@@ 5 Result
This section provides a detailed overview of the experiments conducted for the annotation of three distinct tasks, utilizing five large language models (LLMs) and six BERT-based models. Table 1 summarizes the performance metrics of the fine-tuned BERT-based models, offering a comparative analysis against the performance of each LLM under both few-shot and zero-shot prompting scenarios. The table facilitates a comprehensive evaluation by highlighting key outcomes, enabling a thorough understanding of how each model performs across the different annotation tasks and prompting methods.
@@ 5.1 Key Findings
Among the LLMs evaluated, GPT-4o achieved the best results compared to others, including Llama 3.1 8B, Gemma 2 (2B and 9B), and Gemini 1.0 pro. However, both open-source and closed-source LLMs exhibited notable limitations in providing accurate and reliable annotations.
While few-shot prompting techniques yielded better accuracy than zero-shot approaches, they still fell short of the performance delivered by BERT-based models. This suggests that, despite the increasing popularity of LLMs, BERT-based models continue to be highly relevant, particularly for Indic languages.
@@ 6 Conclusion
Our study demonstrates that while LLMs like GPT, Gemini, Gemma, and Llama show potential, they currently fall short of being reliable annotators for low-resource languages like Marathi, particularly for complex tasks. BERT-based models continue to outperform LLMs in these contexts, suggesting they remain essential for accurate annotation in Indic languages. These findings indicate that further advancements are required in LLMs to make them viable alternatives for human annotations. Additionally, this research highlights the need for developing more robust models tailored to the specific nuances of low-resource languages to reduce dependence on human annotators.
7 Acknowledgement
This work was carried out under the mentorship of L3Cube, Pune. We would like to express our gratitude towards our mentor, for his continuous support and encouragement. This work is a part of the L3Cube-MahaNLP project Joshi (2022).

### v2

@@ Abstract
Low-resource languages face significant challenges due to the lack of sufficient linguistic data, resources, and tools for tasks such as supervised learning, annotation, and classification. This shortage hinders the development of accurate models and datasets, making it difficult to perform critical NLP tasks like sentiment analysis or hate speech detection. To bridge this gap, Large Language Models (LLMs) present an opportunity for potential annotators, capable of generating datasets and resources for these underrepresented languages. In this paper, we focus on Marathi, a low-resource language, and evaluate the performance of both closed-source and open-source LLMs as annotators, while also comparing these results with fine-tuned BERT models. We assess models such as GPT-4o and Gemini 1.0 Pro, Gemma 2 (2B and 9B), and Llama 3.1 (8B and 405B) on classification tasks including sentiment analysis, news classification, and hate speech detection. Our findings reveal that while LLMs excel in annotation tasks for high-resource languages like English, they still fall short when applied to Marathi. Even advanced models like GPT-4o and Llama 3.1 405B underperform compared to fine-tuned BERT-based baselines, with GPT-4o and Llama 3.1 405B trailing fine-tuned BERT by accuracy margins of 10.2% and 14.1%, respectively. This highlights the limitations of LLMs as annotators for low-resource languages.
@@ 1 Introduction
India is a multilingual nation with various regional languages and most of these languages fall under the low-resource (LR) category. Low resource languages are languages such as Marathi and Hindi that lack annotated training datasets and have very few task-specific resources compared to high resource languages such as Spanish and English.
@@
Specifically, we evaluated models such as GPT-4o, Gemini 1.0 Pro, Gemma 2 (2B and 9B), Llama 3.1 (8B and 405B) across multiple tasks, including 3-class sentiment analysis, 2-class, and 4-class hate speech detection, as well as news classification based on headlines, long paragraphs, and full documents.
The key contributions of this research work are as follows:
We have conducted a first-of-its-kind detailed comparative study between fine-tuned BERT models and large language models (LLMs), by evaluating their potential to be used as annotators for a low-resource language, Marathi.
We observe that the average results of the Few-shot prompting technique outperform the average result of the Zero-shot prompting technique in all the models tested.
We have provided valuable insights into the effectiveness of both open- and closed-source large language models (LLMs), including GPT-4o, Llama 3.1 405B, Llama 3.1 8B, Gemma 2 9B, Gemma 2 2B, and Gemini 1.0 Pro, on tasks such as Marathi Sentiment Analysis, Hate Speech Detection, and News Categorization. Our results strongly demonstrate that LLMs are still not fully reliable for annotation tasks in Indic languages.
Model ranking, based on accuracy metrics, is GPT-4o > Llama-3.1-405B > Gemini 1.0 Pro > Gemma 2 9B > Llama 3.1 8B > Gemma-2-2B.
The paper is structured as follows: Section 2 provides a concise review of prior research on data annotation and the use of LLMs. In Section 3, we detail the datasets used and the Section 4 describes models employed in our evaluation. Section 5 describes the experimental setup and the APIs leveraged to evaluate the LLMs. Section 6 presents the results, along with a comparative analysis of various LLMs and BERT-based models, highlighting the key findings of our research. Finally, in Section 7, we conclude our discussion.
@@ 2 Literature Review
Bias is another concern with LLMs. Bavaresco et al. (2024) introduced JUDGE-BENCH to evaluate LLM biases, noting that training data heavily influences model outputs, which can be problematic in annotating complex or sensitive tasks in low-resource languages. While LLMs used for high-resource languages (HRL) are giving promising results, that is not the case for low-resource languages. Nasution and Onan (2024) explored ChatGPT-4’s performance in annotation tasks across languages like Turkish and Indonesian, offering insights into LLM applicability for Low Resource Language(LRL), a relevant consideration for our focus on Marathi.
Dataset | Tech | Llama 3.1 8B | Gemma 2 2B | Gemma 2 9B | Gemini 1.0 Pro | Llama 3.1 405B | GPT-4o | Fine Tuned BERT |
MahaSent | ZS | 0.76 | 0.71 | 0.69 | 0.78 | 0.77 | 0.79 | 0.80 |
FS | 0.79 | 0.76 | 0.78 | 0.76 | 0.81 | 0.82 | ||
MahaHate-2C | ZS | 0.64 | 0.71 | 0.78 | 0.74 | 0.77 | 0.80 | 0.91 |
FS | 0.78 | 0.72 | 0.82 | 0.72 | 0.82 | 0.82 | ||
MahaHate-4C | ZS | 0.40 | 0.39 | 0.43 | 0.43 | 0.49 | 0.58 | 0.73 |
FS | 0.48 | 0.41 | 0.46 | 0.45 | 0.52 | 0.60 | ||
MahaNews-SHC | ZS | 0.60 | 0.54 | 0.68 | 0.68 | 0.75 | 0.78 | 0.85 |
FS | 0.66 | 0.54 | 0.68 | 0.70 | 0.74 | 0.78 | ||
MahaNews-LPC | ZS | 0.66 | 0.55 | 0.71 | 0.72 | 0.76 | 0.77 | 0.89 |
FS | 0.67 | 0.50 | 0.72 | 0.74 | 0.76 | 0.75 | ||
MahaNews-LDC | ZS | 0.69 | 0.62 | 0.78 | 0.74 | 0.76 | 0.81 | 0.96 |
FS | 0.69 | 0.62 | 0.80 | 0.75 | 0.78 | 0.81 | ||
Average | ZS | 0.625 | 0.587 | 0.678 | 0.682 | 0.716 | 0.755 | 0.857 |
FS | 0.678 | 0.592 | 0.710 | 0.687 | 0.738 | 0.763 |
3 Dataset
In this research, we focus on three major task categories using relevant Marathi datasets:
1) MahaSent Kulkarni et al. (2021); Pingle et al. (2023) – classifies sentiment of Marathi tweets into three classes of positive, negative, or neutral categories.
2) MahaHate Patil et al. (2022) – measures the level of abusive and hostile content in Marathi text. This dataset includes two supervised tasks: MahaHate 2-Class, which categorizes content as either HATE or NOT, and MahaHate 4-Class, which provides finer distinctions with categories: Hate (HATE), Offensive (OFFN), Profane (PRFN), and Not (NOT).
3) MahaNews Mittal et al. (2023); Mirashi et al. (2024) – classifies headlines and articles from Marathi news. It comprises three supervised datasets: Short Headlines Classification (SHC), Long Document Classification (LDC), and Long Paragraph Classification (LPC), each categorizing news content into 12 classes: Auto, Bhakti, Crime, Education, Fashion, Health, International, Manoranjan, Politics, Sports, Tech, and Travel. The distribution of all the mentioned datasets is provided in Table 2.
Split | MahaSent | MahaHate 2-C | MahaHate 4-C | SHC | LDC | LPC |
Train | 12114 | 30000 | 21500 | 22014 | 22014 | 42870 |
Valid | 1500 | 3750 | 2000 | 2750 | 2750 | 5366 |
Test | 2500 | 3750 | 1500 | 2761 | 2761 | 5357 |
4 Methodolgy
We investigated the distinctions between LLM-generated and human-generated annotations for the Indic language, Marathi, using a comparative methodology, and analyzed the results with fine-tuned BERT-based models for detailed insights.
@@ 4.1 LLMs
In our annotation experiments, we evaluated the performance of LLMs for the Marathi language using two prompting techniques: zero-shot and few-shot learning. We tested both open-source models (Llama 3.1 8B, Llama 3.1 405B, Gemma 2 2B, and Gemma 2 9B) and closed-source models (Gemini 1.0 Pro, GPT-4o), and compared their results with BERT-based models. The performance of each LLM under both prompting strategies is summarized in Table 1.
@@ 4.2 BERT Based Models
We used fine-tuned BERT-based models to compare performance with LLMs, where MahaSent-MD MahaHate-BERT, MahaNews-SHC-BERT, MahaNews-LPC-BERT, and MahaNews-LDC-BERT are fine-tuned versions of MahaBERT, while MahaHate-multi-RoBERTa has MahaRoBERTa as the base model. Each of these models was fine-tuned on the corresponding datasets, and their respective performances are summarized in Table 1.
@@ 6 Result
This section provides a detailed overview of the experiments conducted for the annotation of three distinct tasks, utilizing six large language models (LLMs) and six BERT-based models (BERT model fine-tuned on target task). Table 1 summarizes the performance metrics of the fine-tuned BERT-based models, offering a comparative analysis against the performance of each LLM under both few-shot and zero-shot prompting scenarios. The table facilitates a comprehensive evaluation by highlighting key outcomes, enabling a thorough understanding of how each model performs across the different annotation tasks and prompting methods.
@@ 6.1 Key Findings
Among the LLMs evaluated, GPT-4o achieved the best results compared to others, including Llama 3.1 8B, Gemma 2 (2B and 9B), and Gemini 1.0 pro. However, both open-source and closed-source LLMs exhibited notable limitations in providing accurate and reliable annotations. Our results also demonstrate that closed LLMs like GPT-4o and Gemini 1.0 Pro outperform open LLMs namely Llama and Gemma 2B and 9B but they still underperform when compared to finetuned BERT for almost all datasets.
Our evaluation ranks the models as GPT-4o > Llama 3.1 405B > Gemini 1.0 Pro > Gemma 2 9B > Llama 3.1 8B > Gemma 2 2B, highlighting that open source Llama 3.1 405B outperforms Gemini 1.0 Pro and is second only to GPT-4o.
Compared to zero-shot prompting, few-shot prompting produced more accurate results because it gave examples of the desired input-output behavior, helping the model to understand the task’s context and expectations better. We observe an absolute increase in the average accuracy of few-shot prompting by 5.3%, 0.5%, 3.2%, 0.5%, 2.2%, and 0.8% compared to zero-shot prompting for models Llama 3.1 8B, Gemma 2 2B, Gemma 2 9B, Gemini 1.0 Pro, Llama 3.1 405B, and GPT-4o, respectively. While few-shot prompting techniques yielded better accuracy than zero-shot approaches, they still fell short of the performance delivered by BERT-based models.
The average accuracy gap between open-source and closed-source models is 6.7%, while the difference between closed-source models and fine-tuned BERT-based models is 13.9%, highlighting the lack of effective LLMs for low-resource languages.
BERT-based fine-tuned models on target task outperformed LLMs in the classification tasks for Marathi language because finetuning enabled better knowledge extraction and alignment with the task’s requirements. On the other hand, LLMs, despite being trained on vast amounts of general data, lack performance on low resource languages and task-specific optimization, which limits their ability to extract the most relevant features for a specific task. This suggests that, despite the increasing popularity of LLMs, BERT-based models continue to be highly relevant, particularly for Indic languages.
We also note that the difference in the results of BERT-based models and the LLMs is comparatively less for easy tasks like the Sentiment Analysis Task, i.e. in the MahaSent dataset. At the same time, the gap is significantly higher for complex tasks like the Hate Classification and News Classification tasks in favor of BERT-based models.
@@ 7 Conclusion
Our study demonstrates that while LLMs like GPT, Gemini, Gemma, and Llama show potential, they currently fall short of being reliable annotators for low-resource languages like Marathi, particularly for complex tasks. BERT-based models continue to outperform LLMs in these contexts, suggesting they remain essential for accurate annotation in Indic languages. These findings indicate that further advancements are required in LLMs to make them viable alternatives for human annotations. This research highlights the need for developing more robust models tailored to the specific details of low-resource languages. This includes the creation of higher-quality, task-specific datasets for low-resource languages, ensuring better representation and reducing biases. Enhanced datasets combined with domain-specific knowledge can significantly improve annotation accuracy.
Acknowledgement
This work was carried out under the mentorship of L3Cube, Pune. We would like to express our gratitude towards our mentor for his continuous support and encouragement. This work is a part of the L3Cube-MahaNLP project Joshi (2022).
Appendix A Appendix
Below are the Few-Shot prompts used for annotation tasks.